# Stability of spherical thin-shell wormholes in scalar-tensor theories

### v1

@@ Abstract
In this article, we construct a family of spherically symmetric thin-shell wormholes within scalar-tensor theories of gravity. In the case of wormholes symmetric across the throat, we study the matter content and analyze the stability of the static configurations under radial perturbations. We apply the formalism to two particular examples involving Einstein-Maxwell gravity coupled to a conformal scalar field. In both scenarios, we show that stable configurations are possible for suitable values of the parameters involved.
@@ 1 Introduction
In the context of the different theories of gravity, there are solutions of the field equations that represent compact objects, such as black holes and wormholes. The existence of wormholes is still hypothetical; they were first introduced by Einstein and Rosen [23] and studied more systematically since the work by Morris and Thorne [24]. These geometries represent bridges that connect two different regions of the Universe or two different universes across a throat [25], where the flare-out condition is fulfilled. A common issue encountered in traversable wormholes is the presence of the so-called exotic matter, i.e. matter that does not satisfy the energy conditions [25], which is not a physically desirable feature. In GR, the exotic matter is always required near the throat [26, 27, 28]; however, its amount can be reduced [29], but at the cost of increasing tensions in it [30]. Many papers concerning wormholes with scalar fields have appeared in the literature, here we can only mention a few [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. On the other hand, a new spacetime can be obtained by the cut and paste of two geometries, which are joined on their boundaries by applying the suitable junction conditions, firstly found in GR [42, 43]. This method is used to construct the denominated thin-shell wormholes [25], that allows to localize and minimize the amount of exotic matter. With this procedure, the new manifold typically has an infinitesimally thin layer of matter at the joining hypersurface, where the throat is placed. The construction, the stability, and other related topics on thin-shell wormholes have been studied within GR [44, 45, 46, 47, 48, 49, 50], in dilaton gravity [51], in Einstein-Gauss-Bonnet [52, 53], in Brans-Dicke [54], in $f(R)$ theory [55, 56, 57, 58, 59, 60], and also in Palatini $f(R)$ gravity [61], among others.
In this paper, we construct thin-shell wormholes with spherical symmetry in Einstein theory non-minimally coupled to a scalar field. In order to have a proper matching at the shell, we apply the junction conditions corresponding to this theory [62]. For a family of wormholes symmetric across the throat, which is located at the shell, we study their dynamical stability and the presence of exotic matter. In particular, for the construction we adopt two black hole solutions within Einstein-Maxwell gravity coupled to a conformally invariant scalar field, one quadratic with a constant scalar field [17] and the other of fourth order with variable scalar field [14]. The structure of this paper is as follows. In Sect. 2, we present the general construction of the wormholes in scalar-tensor gravity. In Sect. 3, we develop the formalism for the stability of wormholes symmetric across the throat. In Sect. 4, we consider two particular examples of spacetimes with a radial electric field, in which we study the stability and the energy conditions. Finally, we summarize our results in Sect. 5. We use units such that $c=1$.
@@ 2 Wormhole construction
In this section, we construct a class of spherically symmetric thin-shell wormholes in scalar-tensor gravity by adopting the so-called junction formalism. We are interested in a class of scalar-tensor theories with a non-minimally coupled real scalar field $\phi$ and a self-interaction potential $U(\phi)$, described by the action in the Jordan frame
@@
where $T_{\mu\nu}$ is the energy-momentum tensor associated with the other matter field Lagrangian $\mathcal{L}_{M}$.
The junction formalism for these theories, introduced in Ref. [62], establishes the conditions for proper matching at a hypersurface $\Sigma$. We denote the jump of any quantity $\Upsilon$ across $\Sigma$ by
@@
$2\varepsilon f(\phi)([K_{\mu\nu}]-h_{\mu\nu}[K])+2Mf^{\prime}(\phi)h_{\mu\nu}=S_{\mu\nu},$ (7)
$M=2\varepsilon f^{\prime}(\phi)[K].$ (8)
In these equations, the prime in $f(\phi)$ represents derivative with respect to $\phi$, $K_{\mu\nu}$ is the second fundamental form, $S_{\mu\nu}$ is the energy-momentum tensor at $\Sigma$, and $M$ is defined by
$M=\varepsilon n^{\mu}[\partial_{\mu}\phi],$ (9)
where $n^{1,2}_{\gamma}$ is the unit normal to $\Sigma$, satisfying $n^{\gamma}n_{\gamma}=\varepsilon$. The hypersurface $\Sigma$ in our case will be timelike and so we set $\varepsilon=1$ from now on (the value $\varepsilon=-1$ corresponds to a spacelike hypersurface). One speaks of a boundary hypersurface when $S_{\mu\nu}\equiv 0$, otherwise of a thin shell of matter.
@@
This cut-and-paste procedure has as a result a new geodesically complete manifold $\mathcal{M}=\mathcal{M}^{1}\cup\mathcal{M}^{2}$. The area $4\pi r^{2}$ is minimal at $r=a$, so the flare-out condition is satisfied there and $\mathcal{M}$ describes a wormhole with two regions connected by the throat $\Sigma$ with radius $a$. We can use the proper radial distance $\ell=\pm\int_{a}^{r}\sqrt{1/A_{1,2}(r)}dr$ in order to introduce a global radial coordinate in $\mathcal{M}$, where the (-) and (++) sign corresponds to $\mathcal{M}^{1}$ and to $\mathcal{M}^{2}$ respectively, with the wormhole throat placed at $\ell=0$.
Apart from the coordinates that describe the 4-dimensional spacetime, we introduce the coordinates $\xi^{i}=(\tau,\theta,\varphi)$ on the surface $\Sigma$, with $\tau$ the proper time on the shell. We let the throat radius $a$ to be a function of $\tau$ in order to study of the stability of the shell under radial perturbations. The line element should be continuous across $\Sigma$ as the time coordinates in each side satisfy that $d\tau^{2}=A_{1}(a)^{2}\left(A_{1}(a)+\dot{a}^{2}\right)^{-1}dt_{1}^{2}=A_{2}(a)^{2}\left(A_{2}(a)+\dot{a}^{2}\right)^{-1}dt_{2}^{2}$, where the dot represents the derivative with respect to $\tau$. The induced metric on $\Sigma$ then reads
@@
The first fundamental form $h_{ij}$ or the metric on the shell associated with the two sides of it is given by
@@
and the second fundamental form or the extrinsic curvature of the shell $K_{ij}$ is calculated from
@@
is the unit normal($n^{\gamma}n_{\gamma}=1$) for each side of the shell. For the metrics given by Eq. (10), they read
@@
$K]=\frac{1}{2}\left(\frac{4\left(\sqrt{A_{1}(r)+\dot{a}^{2}}+\sqrt{A_{2}(r)+\dot{a}^{2}}\right)}{a}+\frac{A_{1}^{\prime}(r)+2\ddot{a}}{\sqrt{A_{1}(r)+\dot{a}^{2}}}+\frac{A_{2}^{\prime}(r)+2\ddot{a}}{\sqrt{A_{2}(r)+\dot{a}^{2}}}\right).$ (20)
By using the definition of $M$ introduced in Eq. (9) and considering that the angular components of the normal vector vanish, we obtain the expression
$M=\frac{\dot{a}^{2}\phi^{\prime}_{2}}{\sqrt{A_{2}+\dot{a}^{2}}}-\frac{\dot{a}^{2}\phi^{\prime}_{1}}{\sqrt{A_{1}+\dot{a}^{2}}}+\sqrt{A_{2}+\dot{a}^{2}}\phi^{\prime}_{2}-\sqrt{A_{1}+\dot{a}^{2}}\phi^{\prime}_{1}.$ (21)
@@
$M+4f^{\prime}(\phi)\left(\frac{\sqrt{A_{1}(r)+\dot{a}^{2}}+\sqrt{A_{2}(r)+\dot{a}^{2}}}{a}+\frac{A_{1}^{\prime}(r)+2\ddot{a}}{4\sqrt{A_{1}(r)+\dot{a}^{2}}}+\frac{A_{2}^{\prime}(r)+2\ddot{a}}{4\sqrt{A_{2}(r)+\dot{a}^{2}}}\right)=0,$ (22)
@@
$\sigma=-2Mf^{\prime}(\phi)-\frac{4f(\phi)}{a}(\sqrt{A_{1}(r)+\dot{a}^{2}}+\sqrt{A_{2}(r)+\dot{a}^{2}})$ (23)
@@
$p=2Mf^{\prime}(\phi)+f(\phi)\left(\frac{2(\sqrt{A_{1}(r)+\dot{a}^{2}}+\sqrt{A_{2}(r)+\dot{a}^{2}})}{a}+\frac{A_{1}^{\prime}(r)+2\ddot{a}}{\sqrt{A_{1}(r)+\dot{a}^{2}}}+\frac{A_{2}^{\prime}(r)+2\ddot{a}}{\sqrt{A_{2}(r)+\dot{a}^{2}}}\right).$ (24)
where $M$ shown in Eq. (21). All the equations above are greatly simplified if the wormhole spacetime is symmetric across the throat, i.e. when the solutions adopted at both sides of the shell are the same.
@@ 3 Stability of wormholes symmetric across the throat
For studying the stability of the static configurations, with a throat radius denoted by $a_{0}$, one usually aims to obtain an expression for $\dot{a}^{2}$ and express it in a form of a master equation
@@
where $V(a)$ can be understood as a potential. The stable solutions under radial perturbations are those that fulfill the condition $V^{\prime\prime}(a_{0})>0$. In our case, it appears quite cumbersome to solve the dynamics in terms of $\dot{a}^{2}$ in a generalized way. However, our study simplifies for thin-shell wormholes symmetric across the throat. In this case, the two spacetimes that are joined at $\Sigma$ are the same, thus $A_{1}(r)=A_{2}(r)=A(r)$ and $\phi_{1}(r)=\phi_{2}(r)=\phi(r)$. This also implies that $\partial_{\mu}\phi]=0$ and consequently $M=0$ because of Eq. (9). The junction condition (8) then reduces to $[K]=0$, which using Eq. (20) takes the form
$8f^{\prime}(\phi)\left(\frac{\sqrt{A(r)+\dot{a}^{2}}}{a}+\frac{A^{\prime}(r)+2\ddot{a}}{4\sqrt{A(r)+\dot{a}^{2}}}\right)=0.$ (26)
Since $f^{\prime}(\phi)\neq 0$ beyond GR, we can rewrite it as
$4(A(r)+\dot{a}^{2})+a(A^{\prime}(r)+2\ddot{a})=0.$ (27)
By using (27) and that $2\ddot{a}=d\dot{a}^{2}/da$, we can find an expression for $\dot{a}^{2}$ in the form of Eq. (25), in which we understand the following expression as the potential
$V(a)=A(a)-\frac{a_{0}^{4}}{a^{4}}A(a_{0}).$ (28)
The second order Taylor expansion of this potential reads
@@
It is not difficult to verify that $V(a_{0})=0$ and with the aid of Eq. (27) also that $V^{\prime}(a_{0})=0$, while the second derivative of the potential evaluated at $a_{0}$ takes the form
$V^{\prime\prime}(a_{0})=-20\frac{A(a_{0})}{a_{0}^{2}}+A^{\prime\prime}(a_{0}).$ (30)
The throat radius $a_{0}$ should satisfy the junction condition (27), which reduces to
$4A(a_{0})+a_{0}A^{\prime}(a_{0})=0.$ (31)
The perfect fluid at the shell has an energy density (23) and a pressure (24) whose expressions in the static case read, respectively

### v2

@@ Abstract
In this article, we construct a family of spherically symmetric thin-shell wormholes within scalar-tensor theories of gravity. In the case of wormholes symmetric across the throat, we study the matter content and analyze the stability of the static configurations under radial perturbations. We apply the formalism to a particular example involving Einstein-Maxwell gravity coupled to a conformally invariant scalar field. We show that stable configurations are possible for suitable values of the parameters involved.
@@ 1 Introduction
In the context of the different theories of gravity, there are solutions of the field equations that represent compact objects, such as black holes and wormholes. The existence of wormholes is still hypothetical; they were first introduced by Einstein and Rosen [23] and studied more systematically since the work by Morris and Thorne [24]. These geometries represent bridges that connect two different regions of the Universe or two different universes across a throat [25], where the flare-out condition is fulfilled. A common issue encountered in traversable wormholes is the presence of the so-called exotic matter, i.e. matter that does not satisfy the energy conditions [25], which is not a physically desirable feature. In GR, the exotic matter is always required near the throat [26, 27, 28]; however, its amount can be reduced [29], but at the cost of increasing tensions in it [30]. Many papers concerning wormholes with scalar fields have appeared in the literature, here we can only mention a few [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. On the other hand, a new spacetime can be obtained by the cut and paste of two geometries, which are joined on their boundaries by applying the suitable junction conditions, firstly found in GR [43, 44]. This method is used to construct the denominated thin-shell wormholes [25], that allows to localize and minimize the amount of exotic matter. With this procedure, the new manifold typically has an infinitesimally thin layer of matter at the joining hypersurface, where the throat is placed. The construction, the stability, and other related topics on thin-shell wormholes have been studied within GR [45, 46, 47, 48, 49, 50, 51], in dilaton gravity [52], in Einstein-Gauss-Bonnet [53, 54], in Brans-Dicke [55], in $f(R)$ theory [56, 57, 58, 59, 60, 61, 62], and also in Palatini $f(R)$ gravity [63], among others.
In this paper, we construct thin-shell wormholes with spherical symmetry in Einstein theory non-minimally coupled to a scalar field. In order to have a proper matching at the shell, we apply the junction conditions corresponding to this theory [64]. For a family of wormholes symmetric across the throat, which is located at the shell, we study their dynamical stability and the presence of exotic matter. In particular, for our construction we adopt a black hole solution within Einstein-Maxwell gravity coupled to a conformally invariant scalar field, having a constant scalar field [17]. The structure of this paper is as follows. In Sect. 2, we present the general construction of the wormholes in scalar-tensor gravity. In Sect. 3, we develop the formalism for the stability of wormholes symmetric across the throat. In Sect. 4, we consider a particular example of an spacetime with a radial electric field, in which we study the stability and the energy conditions. Finally, we summarize our results in Sect. 5. We use units such that $c=1$.
@@ 2 Wormhole construction
In this section, we construct a class of spherically symmetric thin-shell wormholes in scalar-tensor gravity by adopting the so-called junction formalism. We are interested in a class of scalar-tensor theories with a non-minimally coupled real scalar field $\phi$, a self-interaction potential $U(\phi)$, and an arbitrary function $f(\phi)$, described by the action in the Jordan frame
@@
where the prime on the functions $f(\phi)$ and $U(\phi)$ represents the derivative with respect to $\phi$, and $T_{\mu\nu}$ is the energy-momentum tensor associated with the other matter field Lagrangian $\mathcal{L}_{M}$.
The junction formalism for these theories, introduced in Ref. [64], establishes the conditions for proper matching of two manifolds $\mathcal{M}^{1}$ and $\mathcal{M}^{2}$ at a hypersurface $\Sigma$. We denote the jump of any quantity $\Upsilon$ across $\Sigma$ by
@@
$2\varepsilon f(\phi)([K_{\mu\nu}]-h_{\mu\nu}[K])+2\Omega f^{\prime}(\phi)h_{\mu\nu}=S_{\mu\nu},$ (7)
$\Omega=2\varepsilon f^{\prime}(\phi)[K].$ (8)
In these equations, $h_{\mu\nu}$ is the first fundamental form, $K_{\mu\nu}$ is the second fundamental form, $K=K^{\mu}_{\;\;\mu}$ is its trace, $S_{\mu\nu}$ is the energy-momentum tensor at $\Sigma$, and $\Omega$ is defined by
$\Omega=\varepsilon[n^{\mu}\partial_{\mu}\phi],$ (9)
where $n^{\gamma}$ is the unit normal to $\Sigma$ that points from $\mathcal{M}^{1}$ to $\mathcal{M}^{2}$, satisfying $n^{\gamma}n_{\gamma}=\varepsilon$. The hypersurface $\Sigma$ in our case will be timelike and so we set $\varepsilon=1$ from now on (the value $\varepsilon=-1$ corresponds to a spacelike hypersurface). One speaks of a boundary hypersurface when $S_{\mu\nu}\equiv 0$, otherwise, of a thin shell of matter.
@@
This cut-and-paste procedure has as a result a new geodesically complete manifold $\mathcal{M}=\mathcal{M}^{1}\cup\mathcal{M}^{2}$. The area $4\pi r^{2}$ is minimal at $r=a$, so the flare-out condition is satisfied there and $\mathcal{M}$ describes a wormhole with two regions connected by the throat $\Sigma$ with radius $a$. We can use the proper radial distance $\ell=\pm\int_{a}^{r}\sqrt{1/A_{1,2}(r)}dr$ in order to introduce a global radial coordinate in $\mathcal{M}$, where the (-) and (++) sign refers to $\mathcal{M}^{1}$ and to $\mathcal{M}^{2}$ respectively, with the wormhole throat placed at $\ell=0$.
Apart from the coordinates that describe the 4-dimensional spacetime, we introduce the coordinates $\xi^{i}=(\tau,\theta,\varphi)$ on the surface $\Sigma$, with $\tau$ the proper time on the shell. We let the throat radius $a$ to be a function of $\tau$ in order to study of the stability of the shell under radial perturbations. The line element should be continuous across $\Sigma$, then the time coordinates at each side satisfy that $d\tau^{2}=A_{1}(a)^{2}\left(A_{1}(a)+\dot{a}^{2}\right)^{-1}dt_{1}^{2}=A_{2}(a)^{2}\left(A_{2}(a)+\dot{a}^{2}\right)^{-1}dt_{2}^{2}$, where the dot represents the derivative with respect to $\tau$. The induced metric on $\Sigma$ then reads
@@
The first fundamental form $h_{ij}$ (the metric on the shell associated with the two sides of it) is given by
@@
and the second fundamental form (or extrinsic curvature) of the shell $K_{ij}$ is calculated from
@@
is the unit normal for each side of the shell. In this equation, the upper sign refers to $\mathcal{M}^{2}$ and the lower one to $\mathcal{M}^{1}$. For the metrics given by Eq. (10), they read
@@
$K]=\frac{2\left(\sqrt{A_{1}(a)+\dot{a}^{2}}+\sqrt{A_{2}(a)+\dot{a}^{2}}\right)}{a}+\frac{A_{1}^{\prime}(a)+2\ddot{a}}{2\sqrt{A_{1}(a)+\dot{a}^{2}}}+\frac{A_{2}^{\prime}(a)+2\ddot{a}}{2\sqrt{A_{2}(a)+\dot{a}^{2}}}.$ (20)
By using the definition of $\Omega$ introduced in Eq. (9) and considering that the angular components of the unit normal vanish, we obtain the expression
$\Omega=\frac{\left(A_{2}(a)+2\dot{a}^{2}\right)\phi^{\prime}_{2}(a)}{\sqrt{A_{2}(a)+\dot{a}^{2}}}+\frac{\left(A_{1}(a)+2\dot{a}^{2}\right)\phi^{\prime}_{1}(a)}{\sqrt{A_{1}(a)+\dot{a}^{2}}}.$ (21)
@@
$\Omega+4f^{\prime}(\phi)\left(\frac{\sqrt{A_{1}(a)+\dot{a}^{2}}+\sqrt{A_{2}(a)+\dot{a}^{2}}}{a}+\frac{A_{1}^{\prime}(a)+2\ddot{a}}{4\sqrt{A_{1}(a)+\dot{a}^{2}}}+\frac{A_{2}^{\prime}(a)+2\ddot{a}}{4\sqrt{A_{2}(a)+\dot{a}^{2}}}\right)=0,$ (22)
@@
$\sigma=-2\Omega f^{\prime}(\phi)-\frac{4f(\phi)}{a}(\sqrt{A_{1}(a)+\dot{a}^{2}}+\sqrt{A_{2}(a)+\dot{a}^{2}})$ (23)
@@
$p=2\Omega f^{\prime}(\phi)+f(\phi)\left(\frac{2(\sqrt{A_{1}(a)+\dot{a}^{2}}+\sqrt{A_{2}(a)+\dot{a}^{2}})}{a}+\frac{A_{1}^{\prime}(a)+2\ddot{a}}{\sqrt{A_{1}(a)+\dot{a}^{2}}}+\frac{A_{2}^{\prime}(a)+2\ddot{a}}{\sqrt{A_{2}(a)+\dot{a}^{2}}}\right),$ (24)
where $\Omega$ is shown in Eq. (21). All the equations above are greatly simplified if the wormhole spacetime is symmetric across the throat, i.e. when the solutions adopted at both sides of the shell are the same.
@@ 3 Stability of wormholes symmetric across the throat
For the study of the stability of the static configurations with a throat radius denoted by $a_{0}$, one usually aims to express $\dot{a}^{2}$ in the form of a master equation
@@
where $V(a)$ can be understood as a potential. The stable solutions under radial perturbations are those that fulfill the condition $V^{\prime\prime}(a_{0})>0$. In our case, it appears quite cumbersome to solve the dynamics in terms of $\dot{a}^{2}$ in a generalized way. However, our study is simplified for thin-shell wormholes symmetric across the throat. In this case, the two spacetimes that are joined at $\Sigma$ are the same, thus $A_{1}(r)=A_{2}(r)=A(r)$, $\phi_{1}(r)=\phi_{2}(r)=\phi(r)$, and also $\phi^{\prime}_{1}(r)=\phi_{2}^{\prime}(r)=\phi^{\prime}(r)$. Consequently, from Eq. (21) we obtain
$\Omega=\frac{2\left(A(a)+2\dot{a}^{2}\right)\phi^{\prime}(a)}{\sqrt{A(a)+\dot{a}^{2}}}.$ (26)
The junction condition (8), which for a spherical shell is shown in Eq. (22), then takes the form
$\Omega+8f^{\prime}(\phi)\left(\frac{\sqrt{A(a)+\dot{a}^{2}}}{a}+\frac{A^{\prime}(a)+2\ddot{a}}{4\sqrt{A(a)+\dot{a}^{2}}}\right)=0.$ (27)
We can rewrite it as
$a\left(A(a)+2\dot{a}^{2}\right)\phi^{\prime}(a)+4f^{\prime}(\phi)\left(A(a)+\dot{a}^{2}\right)+af^{\prime}(\phi)\left(A^{\prime}(a)+2\ddot{a}\right)=0,$ (28)
or reordering terms
$2af^{\prime}(\phi)\ddot{a}+\left(4f^{\prime}(\phi)-2a\phi^{\prime}(a)\right)\dot{a}^{2}+af^{\prime}(\phi)A^{\prime}(a)+\left(4f^{\prime}(\phi)-a\phi^{\prime}(a)\right)A(a)=0.$ (29)
We look for an expression of $\dot{a}^{2}$ in the form of Eq. (25). With the assumption $f^{\prime}(\phi)\neq 0$, the help of the definitions
$z(a)=\dot{a}^{2}+A(a),\qquad y(a)=\frac{4}{a}-\frac{2\phi^{\prime}(a)}{f^{\prime}(\phi)},\qquad x(a)=\frac{A(a)\phi^{\prime}(a)}{f^{\prime}(\phi)},$ (30)
and using that $2\ddot{a}=d\dot{a}^{2}/da$, we can rewrite Eq. (29) in the form
$z^{\prime}(a)+y(a)z(a)+x(a)=0,$ (31)
which is a first order ordinary linear equation with non-constant coefficients. For regular enough functions $x(a)$ and $y(a)$ this equation can be formally solved to give $z(a)$, and consequently it allows to obtain $\dot{a}^{2}=z(a)-A(a)\equiv-V(a)$. But from a practical point of view, in the particular cases, as $x(a)$ and $y(a)$ depend on the metric function $A(a)$ and also on the functions $f(\phi)$ and $\phi(a)$ associated to the scalar field, finding the expression of the potential and its first and second derivatives can be a rather difficult task.
An interesting situation where the potential can be readily calculated is when $\phi^{\prime}(a)=0$, for which we have that $n^{\mu}\partial_{\mu}\phi]=0$ and then $[K]=0$. In this case, from Eq. (28) the throat radius should fulfill
$4f^{\prime}(\phi)\left(A(a)+\dot{a}^{2}\right)+af^{\prime}(\phi)\left(A^{\prime}(a)+2\ddot{a}\right)=0,$ (32)
which for the static configuration reduces to
$4A(a_{0})+a_{0}A^{\prime}(a_{0})=0$ (33)
when $f^{\prime}(\phi)\neq 0$. The differential equation (31) takes the simple form
$z^{\prime}(a)+\frac{4}{a}z(a)=0,$ (34)
having the solution $z(a)=C/a^{4}$, with $C$ a constant. The potential then reads
$V(a)=A(a)-\frac{a_{0}^{4}}{a^{4}}A(a_{0}),$ (35)
where the constant $C$ has been obtained from the condition $V(a_{0})=0$. The second order Taylor expansion of this potential takes the form
@@
Besides from $V(a_{0})=0$, with the aid of Eq. (33) it is not difficult to verify that $V^{\prime}(a_{0})=0$, while the second derivative of the potential evaluated at $a_{0}$ reads
$V^{\prime\prime}(a_{0})=A^{\prime\prime}(a_{0})-20\frac{A(a_{0})}{a_{0}^{2}}.$ (37)
The stable solutions satisfy that $V^{\prime\prime}(a_{0})>0$.
The perfect fluid at the shell has an energy density (23) and a pressure (24), whose expressions in the static case read, respectively